# InvertAvatar: Incremental GAN Inversion for Generalized Head Avatars

### v2

@@ 3.3 Incremental GAN inversion
We transition from a one-shot to a multi-frame inversion structure using temporal aggregation networks $E_{tex\_rec}$ and $E_{tri\_rec}$ equipped with recurrent decoders. Given an input video clip $\mathbf{I}^{N}={\left\{I_{t}\right\}}_{0}^{N}$, we aim at fusing the information of all frames to generate canonical representation $F_{tex}$ and $F_{tri}$. First, the first frame is input to the $E_{latent}$ for $W+$ space latent code $w$ and initial canonical features $F_{tex_{w}}$ and $F_{tri_{w}}$. According to the poses and expressions of the input sequence, we synthesize a fake image sequence and calculate the residual map for each frame $\Delta{I_{t}}=I_{t}-\hat{I_{t}}$. We sequentially input the observation and residual of each frame into the network $E_{tex\_rec}$ and $E_{tri\_tex}$ respectively on the UV domain and the picture domain.
Performing in a seq2one fashion, the recurrent decoders of $E_{tex\_rec}$ and $E_{tri\_tex}$ merge frame-wise feature maps in the temporal domain to predict one $F_{tri-SFT}$ and $F_{tex_{off}}$. Specifically, the recurrent decoder adopts ConvGRU [5] at each scale to aggregate temporal information, which is defined as:
$\begin{split}&\ z_{t},r_{t}=\sigma(Conv(f_{t},h_{t-1}))\\ &\ o_{t}=tanh(Conv(f_{t},r_{t}*h_{t-1}))\\ &\ h_{t}=z_{t}*h_{t-1}+(1-z_{t})*o_{t}\end{split}$ (1)
where operator *** denotes element-wise product. For each ConvGRU block, the feature map of one frame $f_{t}$ is input to calculate the update weights map $z_{t}$ and reset weights map $r_{t}$, to merge the current frame’s information into hidden state $h_{t}$. With the initial recurrent state $h_{0}$ all zero tensor, after inputting all frames in turn, the final fully updated hidden state serves as the output.
@@
Table 1: Quantitative Evaluation for one-shot avatar reconstruction and animation.
Figure 3: Qualitative comparison on the CelebA-HQ for one-shot avatar reconstruction. The results demonstrate the superior performance of our method in terms of realistic appearance recovery and fine-grained expression control.
4 Experiments
4.1 Experiment Setting
@@
We mainly compare our method with image-driven-based generalized head avatar reenactment methods, including a 2D face reenactment method [55] and two 3D avatar synthesis methods [31, 24]. Besides, we also compare with an optimization-based baseline Next3D-PTI, which incorporates an animatable 3D GAN [40] and pivotal tuning inversion [37].
@@
We evaluate reconstruction and reenactment quality using peak signal-to-noise ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) [58] for synthetic quality, cosine similarity (CSIM) [10] for identity preservation, and average keypoint distance (AKD) using a SOTA off-the-shelf facial landmarks detector [^1] https://github.com/google/mediapipe for reenactment quality.
@@
Table 2: Quantitative Evaluation on VFHQ-Test using multiple source images.
Figure 4: Qualitative comparison on the VFHQ-Test for few-shot avatar reconstruction and animation. Our method integrates multi-frame observations to improve shape reconstruction and texture recovery, as well as retaining high fidelity.
4.2 Comparisons
4.2.1 One-shot Avatar Reconstruction
@@
We evaluate same-identity and cross-identity reenactment using various methods. For self-reenactment, the first frame of HDTF test videos served as the source, with the next 500 frames as the driving video. In cross-identity cases, the first 200 frames of one HDTF clip compose the driving video, and the first frame of other videos as source images. Besides, following [55], We also transfer motions from videos to CelebA-HQ monocular images [22], using test videos from HDTF and VFHQ, creating 100,000 synthesized images from 1000 CelebA-HQ images and 25 video clips  [^2] In addition to 19 videos from the testing split of HDTF, we also select 6 videos with rich expression changes from VFHQ as supplements.. Without ground truth for cross-reenactment, we relied on CSIM and AKD metrics for evaluation.
@@
Tab.1 presents quantitative evaluations on the HDTF and CelebA-HQ datasets, while Fig. 3 showcases the qualitative results of cross-identity reenactment on the CelebA-HQ dataset. Our method shows superior generalization on these unseen datasets, surpassing others in identity preservation, appearance recovery, and expression control. It outperforms baselines in capturing fine details such as hair strands and earrings. Additionally, it also excels in expression transfer, offering precise control over nuanced expressions like asymmetrical mouth movements and eye gazes. In comparison, ROME [24] fails to capture high-frequency details, StyleHeat [55] struggles with inner mouth synthesis, and OTAvatar [31] is unable to generate photo-realistic facial avatars with compact latent vector encoding. Although Next3D-PTI can reconstruct avatars with test-time optimization, it falls short in precise expression control.
4.2.2 Few-shot Avatar Reconstruction
@@
Without readily available benchmarks for few-shot avatar reconstruction, we adapt one-shot methods, both our method and OTAvatar [31], to accommodate multiple source images. This adaptation involves reconstructing canonical features from each source image, followed by averaging these features to derive a canonical space representation. Specifically, for OTAvatar, we compute the mean identity latent code. For our one-shot method, we averaged the feature maps $F_{tex}$ and $F_{tri}$ in both two canonical spaces.
@@
Due to the diverse expressions and poses in VFHQ [53], we evaluate using a test dataset of 100 VFHQ videos, each with 300 frames, excluding training set subjects. We sample input frames evenly from the initial 200 frames and use the last 100 frames for performance evaluation.
@@
For efficiency, OTAvatar[31] and Next3D-PTI require lengthy optimization for novel identities. In contrast, our learning-based solution achieves rapid avatar reconstruction within one second, with the time cost increasing only marginally with the number of input frames.
4.3 Ablation Studies
4.3.1 Recurrent mechanism for temporal aggregation
@@
Figure 5: With a continuous input video stream, our method incrementally refines facial shape and texture details, in contrast to the fixed-window baseline “ConvFusion_avg”, which tends to produce blurry outcomes.
Figure 6: LPIPS over the number of input frames on VFHQ-test. With longer sequences, our method shows improving and converging metrics, affirming its proficiency in long-term temporal aggregation, unlike fixed-window baselines that degrade as source image count consistently increases.
4.3.2 Encoder design for canonical space inversion
@@
We set a baseline with posed images as inputs to $E_{tex}$. As shown in Fig. 7 and Tab. 3, our neural texture encoder, designed to prioritize pixel-aligned image-to-image translation on the UV parameterization, is vital for faithful facial detail recovery.
@@
Table 3: Ablation studies on CelebA-HQ for portrait reconstruction, to validate the effectiveness of our encoder design.
Figure 7: Qualitative ablation study results on the encoder design for canonical space inversion. Zoom in for the best view.
5 Discussion and Conclusion
Limitation. Dependent on a parametric model to control facial expressions, our method struggles with extreme expressions beyond the model’s capability, like frowning and sticking tongue out. Additionally, synthesizing stable, clear dynamic lower teeth remains challenging. Future advancements in more robust facial models capable of tracking these aspects are likely to enhance the performance.
@@ 5 Discussion and Conclusion
Supplementary Material
6 Implementation Details
@@
6.1 Training Animatable 3D GAN
We implement our 3D GAN framework on top of the official PyTorch implementation of Next3D [40], and we illustrate the differences in network architecture between ours and Next3D in Fig 8. Specifically, we rasterize neural texture feature maps at the resolutions of ($32,64,128$) for multi-scale conditions on $G_{face}$.
We adopt many hyperparameters and training strategies of EG3D [8] and Next3D [40] including blurred real images at the beginning, pose conditioned generator, density regularization, learning rates, and two-stage training. As for the deformation-aware discriminator, we replace the input synthetic rendering image with the landmarks counter map to force the generator to focus on the expression changes in the mouth and eyes region. During the training process, we follow Next3D and adopt the same loss terms. Specifically, we train the network on FFHQ [23], using the non-saturating GAN loss with R1 regularization and the density regularization proposed in EG3D. Based on the pretrained model of EG3D, we train our model at a neural rendering resolution of 64 on four 3090 GPUs with a batch size of 16 for roughly 2 days and gradually step the resolution up to 128 for additional 2 days.
Figure 8: Comparison of the network architecture of Next3D against ours. Our more memory-friendly and compact representation allows for efficient training at a neural rendering resolution of $128^{2}$, which significantly improves view consistency.
6.2 Training Encoders
@@
6.2.1 Training $E_{latent}$
For training the latent encoder $E_{latent}$, we employ pixel-wise $L_{1}$ loss, perceptual loss $L_{lpips}$ [58], and identity similarity loss $L_{ID}$ [10]. Besides, we follow [45] and set a latent discriminator $D_{W}$ to discriminate between real samples from the $W$ space and the latent codes predicted by the encoder. The encoder is trained in an adversarial scheme and the total loss can be written as :
$L_{stage1}=L_{1}+\lambda_{lpips}L_{lpips}+\lambda_{id}L_{ID}+L_{adv}^{D_{W}}+L_{adv}^{E_{latent}}$ (2)
@@ 6.2.2 Training $E_{tri}$ and $E_{tex}$
Inspired by [46], leveraging synthetic data, we supervise the training with intermediate feature maps, including neural texture maps $F_{tex}$, tri-plane feature maps $F_{tri}$ and neural rendering feature maps $I_{raw}$. Loss can be written as:
$\begin{split}L_{stage2}=&L_{1}+\lambda_{lpips}L_{lpips}+\lambda_{tri}L_{tri}+\lambda_{tex}L_{tex}\\ &+\lambda_{raw}L_{raw}+\lambda_{adv}L_{adv}\end{split}$ (3)
where $\lambda_{lpips}=1.0$, $\lambda_{tri}=0.001$, $\lambda_{tex}=0.001$, $\lambda_{raw}=1.0$, $\lambda_{adv}=0.1$. $L_{tri}$, $L_{tex}$, $L_{raw}$ are the L1 losses that intuitively compare those quantities synthesized by our generative model and those created by our encoders. $L_{adv}$ is the adversarial loss using a dual discriminator [8], which is trained to differentiate between synthetic images sampled from GAN prior and the corresponding image reconstructed by the encoders. We remove the pose condition for the dual discriminator. On four 3090 GPUs with a batch size of 16, we first train both image encoders without the generative adversarial objective for about 4 days, and then add adversarial loss to finetune encoders for an additional day. We use a learning rate of 1e-4 for encoders and 1e-3 for the discriminator.
6.2.3 Training $E_{tri\_rec}$ and $E_{tex\_rec}$
Building upon pretrained $E_{tri}$ and $E_{tex}$, $E_{tri\_rec}$ and $E_{tex\_rec}$ maintain fixed encoding backbone and only finetune their recurrent decoders. To learn to aggregate temporal information from monocular videos, we leverage 7500 real-world video clips from the VFHQ video dataset [53], each comprising 100 frames. However, since the image quality in the video dataset typically lags behind that of the GAN prior from the FFHQ dataset, training only with the VFHQ data causes blurry and degraded renderings. Hence during training, we simultaneously use the synthetic video data, which has random identity sampled from the GAN latent space, with poses and expressions sampled from the VFHQ video dataset. This additional synthetic data supplementation improves the quality of the generated images.

### v3

@@ 3.4. Incremental GAN inversion
We transition from a one-shot to a multi-frame inversion structure using temporal aggregation networks $E_{tex\_rec}$ and $E_{tri\_rec}$ equipped with recurrent decoders. Given an input video clip $\mathbf{I}^{N}={\left\{I_{t}\right\}}_{0}^{N}$, we aim at fusing the information of all frames to generate canonical representation $F_{tex}$ and $F_{tri}$. First, the first frame is input to the $E_{latent}$ for $W+$ space latent code $w$ and initial canonical features $F_{tex_{w}}$ and $F_{tri_{w}}$. According to the poses and expressions of the input sequence, we synthesize a fake image sequence and calculate the residual map for each frame $\Delta{I_{t}}=I_{t}-\hat{I_{t}}$. We sequentially input the observation and residual of each frame into the network $E_{tex\_rec}$ and $E_{tri\_rec}$ respectively on the UV domain and the picture domain.
Performing in a seq2one fashion, the recurrent decoders of $E_{tex\_rec}$ and $E_{tri\_tex}$ merge frame-wise feature maps in the temporal domain to predict one $F_{tri-SFT}$ and $F_{tex_{off}}$. Specifically, the recurrent decoder adopts ConvGRU (Ballas et al., 2015) at each scale to aggregate temporal information, which is defined as:
(1) $\begin{split}&\ z_{t},r_{t}=\sigma(Conv(f_{t},h_{t-1}))\\ &\ o_{t}=tanh(Conv(f_{t},r_{t}*h_{t-1}))\\ &\ h_{t}=z_{t}*h_{t-1}+(1-z_{t})*o_{t}\end{split}$
where operator * denotes element-wise product. For each ConvGRU block, the feature map of one frame $f_{t}$ is input to calculate the update weights map $z_{t}$ and reset weights map $r_{t}$, to merge the current frame’s information into hidden state $h_{t}$. With the initial recurrent state $h_{0}$ all zero tensor, after inputting all frames in turn, the final fully updated hidden state serves as the output.
@@
Figure 4. Compare our animatable 3D generative model with Next3D (Sun et al., 2023). We extract frames from a driving video clip and use the estimated facial model parameters to animate sampled random virtual avatars. Please zoom in and also refer to our video for more clear comparisons.
4. Experiments
4.1. Experiment Setting
Table 1. Quantitative comparison with Next3D on image synthesis quality, animation accuracy, identity preservation and memory consumption.
[…]
Ours 4.1 0.057 0.853 84.21M
Next3D 3.9 0.248 0.841 164.81M
Table 2. Quantitative Evaluation for one-shot avatar reconstruction and animation.
@@ 4.1. Experiment Setting
Firstly, we compare the utilized animatable 3D GAN prior with Next3D (Sun et al., 2023). Then, we mainly compare our method with image-driven-based generalized head avatar reenactment methods, including a 2D face reenactment method (Yin et al., 2022) and two 3D avatar synthesis methods (Ma et al., 2023; Khakhulin et al., 2022). We also compare with an optimization-based baseline Next3D-PTI, which incorporates an animatable 3D GAN (Sun et al., 2023) and pivotal tuning inversion (Roich et al., 2021). We do not make comparisons with person-specific head avatar methods, since these multi-shot methods cannot generalize to different identities.
@@
We evaluate generation, reconstruction and reenactment quality using Frechet Inception Distance (FID) (Heusel et al., 2017), peak signal-to-noise ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) for synthetic quality, cosine similarity (CSIM) (Deng et al., 2019) for identity preservation, and average keypoint distance (AKD) using a SOTA off-the-shelf facial landmarks detector [^3] https://github.com/google/mediapipe for reenactment quality.
Table 3. Quantitative Evaluation on VFHQ-Test using multiple source images.
@@
4.2. Comparisons
4.2.1. Animatable 3D GAN prior
To validate the superiority of our compact and expressive animatable 3D GAN, we compare our generative model with Next3D. Fig. 4 and Tab. 1 demonstrate qualitative and quantitative results. We measure image quality with FID between the entire FFHQ dataset and 50k generated images using randomly sampled latent codes, camera poses and estimated facial model parameters. Then we randomly sample 100 identities and animate each with randomly sampled 100 images from FFHQ, synthesizing 10000 images. AKD is calculated between the driving images and the animated generated ones, and CSIM is calculated for each identity. The comparison proves that, with roughly half the number of parameters, our generative model achieves the similar synthesis quality as Next3D and significantly improves animation accuracy. In the supplemental video, the video animation results present that Next3D suffers from shaking facial appearance and texture sticking artifacts, while our method achieves better temporal- and view-consistency.
4.2.2. One-shot Avatar Reconstruction
@@
We evaluate same-identity and cross-identity reenactment using various methods. For self-reenactment, the first frame of HDTF test videos served as the source, with the next 500 frames as the driving video. In cross-identity cases, the first 200 frames of one HDTF clip compose the driving video, and the first frame of other videos as source images. Besides, following (Yin et al., 2022), We also transfer motions from videos to CelebA-HQ monocular images (Karras et al., 2018), using test videos from HDTF and VFHQ, creating 100,000 synthesized images from 1000 CelebA-HQ images and 25 video clips  [^4] In addition to 19 videos from the testing split of HDTF, we also select 6 videos with rich expression changes from VFHQ as supplements.. Without ground truth for cross-reenactment, we relied on CSIM and AKD metrics for evaluation.
@@
Tab. 2 presents quantitative evaluations on the HDTF and CelebA-HQ datasets, while Fig. 8 showcases the qualitative results of cross-identity reenactment on the CelebA-HQ dataset. Our method shows superior generalization on these unseen datasets, surpassing others in identity preservation, appearance recovery, and expression control. It outperforms baselines in capturing fine details such as hair strands and earrings. Additionally, it also excels in expression transfer, offering precise control over nuanced expressions like asymmetrical mouth movements and eye gazes. In comparison, ROME (Khakhulin et al., 2022) fails to capture high-frequency details, StyleHeat (Yin et al., 2022) struggles with inner mouth synthesis, and OTAvatar (Ma et al., 2023) is unable to generate photo-realistic facial avatars with compact latent vector encoding. Although Next3D-PTI can reconstruct avatars with test-time optimization, it falls short in precise expression control.
4.2.3. Few-shot Avatar Reconstruction
@@
Without readily available benchmarks for few-shot avatar reconstruction, we adapt one-shot methods, both our method and OTAvatar (Ma et al., 2023), to accommodate multiple source images. This adaptation involves reconstructing canonical features from each source image, followed by averaging these features to derive a canonical space representation. Specifically, for OTAvatar, we compute the mean identity latent code. For our one-shot method, we averaged the feature maps $F_{tex}$ and $F_{tri}$ in both two canonical spaces.
@@
Due to the diverse expressions and poses in VFHQ (Xie et al., 2022), we evaluate using a test dataset of 100 VFHQ videos, each with 300 frames, excluding training set subjects. We sample input frames evenly from the initial 200 frames and use the last 100 frames for performance evaluation.
@@
For efficiency, OTAvatar(Ma et al., 2023) and Next3D-PTI require lengthy optimization for novel identities. In contrast, our learning-based solution achieves rapid avatar reconstruction within one second, with the time cost increasing only marginally with the number of input frames.
4.3. Ablation Studies
4.3.1. Recurrent mechanism for temporal aggregation
@@
Figure 5. With a continuous input video stream, our method incrementally refines facial shape and texture details, in contrast to the fixed-window baseline “ConvFusion_avg”, which tends to produce blurry outcomes.
Figure 6. LPIPS over the number of input frames on VFHQ-test. With longer sequences, our method shows improving and converging metrics, affirming its proficiency in long-term temporal aggregation, unlike fixed-window baselines that degrade as source image count consistently increases.
4.3.2. Encoder design for canonical space inversion
Remove the second stage inversion
We set a baseline only utilizing $E_{latent}$. As shown in Fig. 7 and Tab. 4, without both $E_{tex}$ and $E_{tri}$, the reconstructed portrait cannot recover the identity.
@@ 4.3.2. Encoder design for canonical space inversion
We set a baseline with posed images as inputs to $E_{tex}$. As shown in Fig. 7 and Tab. 4, our neural texture encoder, designed to prioritize pixel-aligned image-to-image translation on the UV domain, is vital for faithful facial detail recovery.
Figure 7. Qualitative ablation study results on the encoder design for canonical space inversion. Zoom in for the best view.
Table 4. Ablation studies on CelebA-HQ for portrait reconstruction, to validate the effectiveness of our encoder design. We sample 1000 images to calculate metrics.
@@
5. Discussion and Conclusion
Limitation. Dependent on a parametric model to control facial expressions, our method is error-prone to tracking errors and struggles with extreme expressions beyond the model’s capability, like frowning and sticking tongue out. Additionally, synthesizing stable, clear dynamic lower teeth remains challenging. Future advancements in more robust facial models capable of tracking these aspects are likely to enhance the performance.
@@ 5. Discussion and Conclusion
Acknowledgements.
This paper is supported by National Key R&\&&D Program of China (2022YFF0902200), the NSFC project No.62125107.
Appendix A Implementation Details
@@
A.1. Animatable 3D GAN
We implement our 3D GAN framework on top of the official PyTorch implementation of Next3D (Sun et al., 2023), and we illustrate the differences in network architecture between ours and Next3D in Fig. 10. Next3D transforms neural texture into orthogonal views and forms tri-plane features. To compensate for the FLAME model’s missing inner mouth area, Next3D utilizes a UNet $G_{teeth}$ to supplement and reintegrate this texture, using another UNet $G_{blending}$ for post-blending. These operations are redundant and lead to heavy parameter size, so we alternately propose a more compact architecture, maintaining the insight of completing inner mouth feature conditioning on global information, but removing heavy networks for specific mouth synthesis. Specifically, we propose a face synthesis module $G_{face}$ which is conditioned on the multi-scale rasterized neural textures. We illustrate its network architecture in Fig. 11. The multi-scale neural textures are rasterized from texture space into tri-plane space and then integrated with $G_{face}$’s corresponding feature maps through alpha blending. Conditioned on such global information, $G_{face}$ is precisely controlled by the deformable mesh model, and gradually inpaint inner mouth features in the forward pass layer by layer. Additionally, our more memory-efficient and compact model design allows for neural rendering resolution of $128^{2}$, significantly enhancing view-consistency. Specifically, we rasterize neural texture feature maps at the resolutions of ($32,64,128$) for multi-scale conditions on $G_{face}$.
A.1.1. Training Details
We adopt many hyperparameters and training strategies of EG3D (Chan et al., 2022) and Next3D (Sun et al., 2023) including blurred real images at the beginning, pose conditioned generator, density regularization, learning rates, and two-stage training. As for the deformation-aware discriminator, we replace the input synthetic rendering image with the landmarks counter map to force the generator to focus on the expression changes in the mouth and eyes region. During the training process, we follow Next3D and adopt the same loss terms. Specifically, we train the network on FFHQ (Karras et al., 2020), using the non-saturating GAN loss with R1 regularization and the density regularization proposed in EG3D. Based on the pretrained model of EG3D, we train our model at a neural rendering resolution of 64 on four 3090 GPUs with a batch size of 16 for roughly 2 days and gradually step the resolution up to 128 for additional 2 days.
Figure 10. Comparison of the network architecture of Next3D against ours. Our more memory-friendly and compact representation allows for efficient training at a neural rendering resolution of $128^{2}$, which significantly improves view consistency.
Figure 11. The detailed architecture of $G_{face}$.
A.2. Training Encoders
@@
A.2.1. Training $E_{latent}$
For training the latent encoder $E_{latent}$, we employ pixel-wise $L_{1}$ loss, perceptual loss $L_{lpips}$ (Zhang et al., 2018), and identity similarity loss $L_{ID}$ (Deng et al., 2019). Besides, we follow (Tov et al., 2021) and set a latent discriminator $D_{W}$ to discriminate between real samples from the $W$ space and the latent codes predicted by the encoder. The encoder is trained in an adversarial scheme and the total loss can be written as :
(2) $L_{stage1}=L_{1}+\lambda_{lpips}L_{lpips}+\lambda_{id}L_{ID}+L_{adv}^{D_{W}}+L_{adv}^{E_{latent}}$
@@ A.2.2. Training $E_{tri}$ and $E_{tex}$
Inspired by (Trevithick et al., 2023), leveraging synthetic data, we supervise the training with intermediate feature maps, including neural texture maps $F_{tex}$, tri-plane feature maps $F_{tri}$ and neural rendering feature maps $I_{raw}$. Loss can be written as:
(3) $\begin{split}L_{stage2}=&L_{1}+\lambda_{lpips}L_{lpips}+\lambda_{tri}L_{tri}+\lambda_{tex}L_{tex}\\ &+\lambda_{raw}L_{raw}+\lambda_{adv}L_{adv}\end{split}$
where $\lambda_{lpips}=1.0$, $\lambda_{tri}=0.001$, $\lambda_{tex}=0.001$, $\lambda_{raw}=1.0$, $\lambda_{adv}=0.1$. $L_{tri}$, $L_{tex}$, $L_{raw}$ are the L1 losses that intuitively compare those quantities synthesized by our generative model and those created by our encoders. $L_{adv}$ is the adversarial loss using a dual discriminator (Chan et al., 2022), which is trained to differentiate between synthetic images sampled from GAN prior and the corresponding image reconstructed by the encoders. We remove the pose condition for the dual discriminator. On four 3090 GPUs with a batch size of 16, we first train both image encoders without the generative adversarial objective for about 4 days, and then add adversarial loss to finetune encoders for an additional day. We use a learning rate of 1e-4 for encoders and 1e-3 for the discriminator.
A.2.3. Training $E_{tri\_rec}$ and $E_{tex\_rec}$
Building upon pretrained $E_{tri}$ and $E_{tex}$, $E_{tri\_rec}$ and $E_{tex\_rec}$ maintain fixed encoding backbone and only finetune their recurrent decoders. To learn to aggregate temporal information from monocular videos, we leverage 7500 real-world video clips from the VFHQ video dataset (Xie et al., 2022), each comprising 100 frames. However, since the image quality in the video dataset typically lags behind that of the GAN prior from the FFHQ dataset, training only with the VFHQ data causes blurry and degraded renderings. Hence during training, we simultaneously use the synthetic video data, which has random identity sampled from the GAN latent space, with poses and expressions sampled from the VFHQ video dataset. This additional synthetic data supplementation improves the quality of the generated images.